OpenStudy (anonymous):

derivative of y=x(9-x)^1/2?

OpenStudy (anonymous):

\[y=x\sqrt{(9-x)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

product rule for this one

OpenStudy (anonymous):

this is what I have: x[-1/x(9-x)^1/2]+(9-x)^1/2(1)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

if so how to simplify

OpenStudy (anonymous):

\[y'=1\times \sqrt{9-x}+x\times \frac{1}{\sqrt{9-x}}\times -1\] and then clean up with some algebra

OpenStudy (anonymous):

really not much to clean up, just \[y'=\sqrt{9-x}-\frac{x}{\sqrt{9-x}}\]

OpenStudy (anonymous):

simplifiedmore?

OpenStudy (anonymous):

or if you want to put it over one denominator, you can

OpenStudy (anonymous):

\[\frac{9-x-x}{\sqrt{9-x}}\] \[\frac{9-2x}{\sqrt{9-x}}\]

OpenStudy (anonymous):

if you want to "simplify" things usually best not to use exponents, but rather radicals

OpenStudy (anonymous):

so√(9-x) *x equals 9-2x right

OpenStudy (anonymous):

on no

OpenStudy (anonymous):

so how did u getthe numerator

OpenStudy (anonymous):

\[y'=\sqrt{9-x}-\frac{x}{\sqrt{9-x}}\]

OpenStudy (anonymous):

is that part ok?

OpenStudy (anonymous):

oh so u subtract

OpenStudy (anonymous):

you want to put i to over the same denominator, so you multiply top and bottom of the first term by \[\sqrt{9-x}\]

OpenStudy (anonymous):

\[y'=\sqrt{9-x}-\frac{x}{\sqrt{9-x}}\] \[y'=\frac{\sqrt{9-x}}{\sqrt{9-x}}\times \sqrt{9-x}-\frac{x}{\sqrt{9-x}}\]

OpenStudy (anonymous):

how come wolfram has: -3(x-6)/2(9-x)^1/2 ?